The general solution u(x,t)=(A sin pcx +B cos pcx )(C sin pct+D cos pct ) with u(0,t)=0 and u(l,t)=0 gives..........
1
Expert's answer
2020-09-29T17:25:28-0400

Given the Dirichlet problem with the initial Dirichlet boundary conditions

"u(0,t)=0=u(L,t)"

Then the problem has at most one solution.

To solve this,

Suppose that the problem has two solutions "u_1" and "u_2" . Setting "u=u_1\u2212u_2" , the function "u" solves


"\begin{cases} 
 u_t-ku_{xx}=0 & x\in [0,L], t>0\\
 u(0,t)=0, u(L,t)=0 & t \geq0 \\
 u(x,0)=0 & x\in [0,L] 
 \end{cases}"

Define the energy


"E(t)=\frac12 \intop_0^L((A\ sin\ pcx+ B\ cos\ pcx)(C\ sin\ pct+ D\ cos\ pct))^2dx"

Differentiating with respect tot, we get

"E'(t)=\frac12 \intop_0^L(C\ sin\ pct+ D\ cos\ pct)(A\ sin\ pcx+ B\ cos\ pcx)(C\ sin\ pct+ D\ cos\ pct)=\intop _0^k u(x,t)u_{xx}(x,t)dx"


Integrating by part and using the boundary conditions, we find that

"E'(t) \intop_0^k u(x,t)v_{xx}(x,t)dx=[ku(x,t)u_x(x,t)]_0^L-\intop_0^Lku_x(x,t)^2dx \leq0"


Hence the energy "E" is decreasing. Since "E(0) = 0" and "E(t)\u22650" , we get that "E\u22610" . his implies that "u(0,t)\u22610" for every "t" . Consequently,"u_1(x,t) =u_2(x,t)" for all "x\u2208[0,L]" and "t\u22650" .